Jim runs 6 kilometers in 50 minutes. At the same rate, how many kilometers would he run in 35 minutes?
Found 2 solutions by stanbon, jim_thompson5910:
Answer by stanbon(75887)

Jim runs 6 kilometers in 50 minutes. At the same rate, how many kilometers would he run in 35 minutes?
---------
Use a proportion::
dist/time = dist/time
x/35 = 6/50
----
x = 35(6/50)
----
x = (7/10)6
----
x = 4.2 km
--------------------
Cheers,
Stan H.
===================

Answer by jim_thompson5910(35256)

let x = number of km ran in 35 min

(# of km)/(# of min) = (# of km)/(# of min)

(6 km)/(50 min) = (x km)/(35 min)

6/50 = x/35

6*35 = 50*x

210 = 50x

50x = 210

x = 210/50

x = 21/5

The answer as a fraction is 21/5

Use a calculator to get 21/5 = 4.2

The answer as a decimal is 4.2 kilometers.